okay i just don't know how to put this as an equation?
Rosa has $3.10 in nickles and dimes. She has 10 fewer nickels than she has dimes. How many dimes and nickels does she have?

Answer by stanbon(57219)

Rosa has $3.10 in nickles and dimes. She has 10 fewer nickels than she has dimes. How many dimes and nickels does she have?
----
Quantity Equation: n = d-10
Value Equation: 10d + 5n = 310 cents
----------------------------------------------
Rearrange and modify the equations:
10d - 10n = 100
10d + 5n = 310
-------------------------
Subtract 1st from 2nd and solve for "n":
15n = 210
n = 14 (# of nickels)
---
Since n = d-10, d = 24 (# of dimes)
========================================
Cheers,
Stan H.
========================================